The perimeter of a rectangle is 200 yards. What are the dimensions of the rectangle if the length is 90 yards more than the width?
Answer by macston(5194)

W=width; L=length=W+90yds; P=perimeter=200yds
P=2(L+W) Substitute for L.
200yds=2(W+90yds+W) Divide each side by 2
100yds=2W+90yds Subtract 90yds from each side.
10yd=2W
5yd=W ANSWER 1: the width is 5 yards.
L=W+90=5+90=95yds ANSWER 2: The length is 95 yards.
CHECK:
P=2(L+W)
200yds=2(95yd+5yd)
200yd=2(100yd)
200yd=200yd